OpenStudy (anonymous):

What is the slope of the line passing through the points (0, −5) and (4, 2)? @ganeshie8 please help mmme

OpenStudy (anonymous):

slope of the line is gives by the formula m=y2-y1/x2-x1

OpenStudy (anonymous):

where y2 =2 y1=-5 x1=0 x2=4

OpenStudy (anonymous):

so wats the anwer?

OpenStudy (anonymous):

s0 m comes out to be 7/4

OpenStudy (anonymous):

What is the slope of the line whose equation is 12 = 4x – 6y? help me with this one @danish071996

OpenStudy (anonymous):

wht is in place of question mark

OpenStudy (anonymous):

nothing its just a question mark lol

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

What is the slope of the line whose equation is 12 = 4x – 6y

OpenStudy (anonymous):

@danish071996

OpenStudy (anonymous):

12=4x+6y

OpenStudy (anonymous):

12 = 4x – 6y

OpenStudy (anonymous):

@ganeshie8 @danish071996

OpenStudy (anonymous):

if its 12=4x+6y then use the formula y=mx+c

OpenStudy (anonymous):

6y=12-4x

OpenStudy (anonymous):

ITS IS EXACTLY HOW I TOLD U NO OTHER WAY

OpenStudy (anonymous):

sorry i m not getting

OpenStudy (anonymous):

and this one @danish071996 What is the slope of the line passing through the points (1, –5) and (4, 1)?

OpenStudy (anonymous):

apply same formula i gave u m=y2-y1/x2-x1

OpenStudy (anonymous):

so wats the anwer -2?

OpenStudy (anonymous):

no its -4/3

OpenStudy (anonymous):

1-5/4-1

OpenStudy (anonymous):

I don't have those as my anwer choices

OpenStudy (anonymous):

what r the choices

OpenStudy (anonymous):

a. 2 b.-4/5 c.5/4 d.-2

OpenStudy (anonymous):

no the answer cn cn be b -4/5 if in the question it is given (-1,5)

OpenStudy (anonymous):

do u know how to do these bcuz they know wat they do

OpenStudy (anonymous):

sorry